Solve the system of equations by the substitution method. (If the system is dependent, enter a general solution in terms of c
{3x + 4y= 12
y= −2x + 2
(x, y) =

Answer by stanbon(75887)

Solve the system of equations by the substitution method. (If the system is dependent, enter a general solution in terms of c
{3x + 4y= 12
y= −2x + 2
-----
Substitute for "y" and solve for "x"::
3x + 4(-2x+2) = 12
3x - 8x + 8 = 12
-5x = 4
x = -4/5
Solve for "y"::
y = -2x+2
y = -2(-4/5)+(10/5)
y = 18/5
----------
Cheers,
Stan H.
-------------